• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 21:48:47 on November 11, 2009

    Entry number 299023

    keyword=Compton tune

    We got the expert of MCC to check the background rates. Closing the slits to have good clean beam at 10 uA yielded a background rate of about 150 Hz/uA at 1650 V. Reza tweaked the injector to find the parameters to improve the background and get the rates low at high current. It seems the photon Compton rates are pretty gang (?) phase of the North Linac, the prebuncher is also crucial at high current ( running without indeed it makes rate really bad ). We made an aperture scan in X and a vertical scan. The signal could be clearly seen at 1600 V with a gain of 1, though the signal seemed to disappear at high current mostly due of increase and background but I wonder if the PMT functions correctly at high current since it draws a significant amount of current. One oddity was we could not see the signal any more when going back to 10 uA so I turned the detector off in case it needed to recover. It seemed to work at 2 uA when the beam was resumed. I asked to Megan to check the linearity of the PMT with the LED with beam to check for any gain change due to high current. It seems the detector sees a lot of low energy background so that might change the gain. So additionnal shielding and fixing the bad vacuum might be the way to go.

    More detailed informations in Arne's http://opweb.acc.jlab.org/CSUEApps/elog02/elog_item.php?elog_id=1498001 Mike Tiefenbach http://opweb.acc.jlab.org/CSUEApps/elog02/elog_item.php?elog_id=1498005 Reza Kazimi http://opweb.acc.jlab.org/CSUEApps/elog02/elog_item.php?elog_id=1497992

    A lot of progress since we could actually see the signal and optimize it but still some work needed since it seems the background increases at 80 uA.